OpenStudy (anonymous):

15. The complement of x ◦ is (90−x) ◦ . always sometimes or never

OpenStudy (mathstudent55):

What are complementary angles?

OpenStudy (anonymous):

two angles that add up to 90 degrees

OpenStudy (mathstudent55):

Great. You have angle x. Let's say its complement is angle y. Since you know their measures add up to 90 degrees, you can write x + y = 90 Right?

OpenStudy (anonymous):

right

OpenStudy (mathstudent55):

Good. We said the complement is y. Solve x + y = 90 for y. What do you get?

OpenStudy (anonymous):

1-89 degrees

OpenStudy (anonymous):

or 45 degrees if both are the same

OpenStudy (mathstudent55):

No. This is how you solve for y. We want y. x is being added to y, so we subtract x from both sides: x + y = 90 -x - x ----------------- y = 90 - x Now we see that the complement, y, is 90 - x. Now compare that with the statement of the problem.

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

I understand

OpenStudy (anonymous):

that would be a never, correct?

OpenStudy (mathstudent55):

The answer is always.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oooh nvm

OpenStudy (anonymous):

I get it now

OpenStudy (anonymous):

thank you!

OpenStudy (mathstudent55):

The problem states that the complement of x is 90 - x We called the complement of x y, and we solved the equation and found out that y is 90 - x, exactly what the problem stated was the complement. That means the answer is always.

OpenStudy (mathstudent55):

You're welcome.